OpenStudy (anonymous):

help plz QAQ !!! An urn contains 20 blue balls 30 yellow balls and 50 red balls.Suppose 5 balls are drawn randomly each at a time without replacement from the urn. let G be the event that exactly one ball is yellow and one ball is red among the 5 chosen balls. (a)Find the probability of Event G. (b)Find the probability of Event G given that at least three chosen balls are blue.

OpenStudy (anonymous):

Okay. I'm a little rusty on probability... But, you first have to add all of the balls together. you have 20+30+50=100, which makes this easier. In a perfect world 3/10 times you will grab a yellow ball, while 5/10it will be red, and the rest (blue) will be 2/10. If you've studied probability this will become a lot easier.

OpenStudy (anonymous):

hmmm i think there is more to it than that

OpenStudy (anonymous):

I agree @satellite73 I just cant remember it all

OpenStudy (anonymous):

your denominator is going to be the number of ways to choose 5 balls out of 100, which is \(\binom{100}{5}\) i would use a calculator

OpenStudy (anonymous):

one ball red, number of ways is \(\binom{50}{1}=50\) one ball yellow \(\binom{30}{1}=30\) and the other 3 are blue, \(\binom{20}{3}=1140\)

OpenStudy (anonymous):

making your final answer \[\frac{50\times 30\times 1140}{\binom{100}{5}}\]

OpenStudy (anonymous):

@satellite73 thank you a lot! How about the question A?